Solve the following equations by the use of the operator D


1. D²y -2Dy + y= sinx +x²


2.D²y -6Dy +6y= e^3x + e^-3x


1
Expert's answer
2021-06-28T16:53:10-0400

1. Let us solve the equation "D^2y -2Dy + y=\sin x +x^2." Its characteristic equation "k^2-2k+1=0" is equivalent to "(k-1)^2=0," and hence has the solutions "k_1=k_2=1." It follows that the general solution is of the form "y=(C_1+C_2x)e^x+y_p," where "y_p=a\sin x+b\cos x+cx^2+dx+e." It follows that "Dy_p=a\cos x-b\sin x+2cx+d" and "D^2y_p=-a\sin x-b\cos x+2c." We have the equality "-a\sin x-b\cos x+2c-2(a\cos x-b\sin x+2cx+d)+a\sin x+b\cos x+cx^2+dx+e=\sin x +x^2."It is equivalent to "-2a\cos x+2b\sin x+cx^2+(d-4c)x+e+2c-2d=\sin x +x^2." It follows that "-2a=0,\ 2b=1,\ c=1,\ d-2c=0,\ e+2c-2d=0." Therefore, "a=0,\ b=\frac{1}{2},\ c=1,\ d=2,\ e=2." We conclude that the general solution is the following:


"y=(C_1+C_2x)e^x+\frac{1}{2}\cos x+x^2+2x+2."



2. Let us solve the equation "D^2y -6Dy +6y= e^{3x} + e^{-3x}." Its characteristic equation "k^2-6k+6=0" is equivalent to "(k-3)^2=3," and hence has the solutions "k_1=3+\sqrt{3}"

and "k_2=3-\sqrt{3}." It follows that the general solution is of the form "y=C_1e^{(3+\sqrt{3})x}+C_2e^{(3-\sqrt{3})x}+y_p," where "y_p=a e^{3x} + be^{-3x}." Therefore, "Dy_p=3a e^{3x} -3be^{-3x}" and "D^2y_p=9a e^{3x} + 9be^{-3x}." We get the equality "9a e^{3x} + 9be^{-3x}-6(3a e^{3x} -3be^{-3x})+6(a e^{3x} +be^{-3x})=e^{3x} +e^{-3x}." It is equivalent to "-3a e^{3x} + 33be^{-3x}=e^{3x} +e^{-3x}." It follows that "-3a=33b=1," and thus "a=-\frac{1}{3},\ b=\frac{1}{33}." We conclude that the general solution is the following:


"y=C_1e^{(3+\sqrt{3})x}+C_2e^{(3-\sqrt{3})x}-\frac{1}{3}e^{3x}+\frac{1}{33}e^{-3x}."